### v1

@@ 4.4 Discussion
In QMCL the surrogate fluxes are computed as the expected value of an observable with respect to a given quantum state. As a result, to compute fluxes that vary strongly over the spatial domain one has to ensure that the used quantum states are very sharp: obtaining a clear maximum in some regions of product space $\Omega$ and being close to zero everywhere else, formally similar to a finite sum of delta Dirac measures. Working with such sharp quantum states has both theoretical and numerical challenges.
@@
Theoretically, producing accurate but very sharp quantum states requires a very precise choice of bandwidth parameter $\epsilon^{\prime}$ used in kernel (9) for bayesian conditioning. Having such a selective kernel means that the resulting quantum state leads to highly accurate fluxes when the correct part of the training dataset is identified as similar to the current resolved state. However, it can lead to strongly inaccurate fluxes when regions of the training dataset are misidentified as similar, compared to using a less selective kernel that leads to wider averaging of the employed quantum observable.
@@
Numerically, working with a very selective kernel can lead to numerical instability, especially when no region of the training dataset is identified as similar to the current observed dynamics. When that happens, the produced feature vector is approximately zero everywhere. During bayesian conditioning, this leads to a quantum state that is also approximately zero everywhere, making it numerically unstable to attempt to normalize that quantum state to unit norm. Working with a wider, less selective kernel makes this scenario less likely to occur, albeit with the tradeoff of less informative conditioning and thus less accurate predicted fluxes.
@@ 5 Conclusion
We have developed a framework for the dynamical closure of spatiotemporal dynamics which combines elements of the quantum mechanical framework developed in [20, 16, 17] and the vector valued spectral analysis (VSA) framework of [22]. We have shown that the present extended framework inherits the symmetry factorization properties of VSA during the basis construction process, and can also be used with training data consisting of multiple dynamical trajectories. Additionally, as with the original QMCL work, the process of embedding classical dynamics into the noncommutative setting of quantum mechanics leads to positivity preserving discretizations of classical observables and probability density functions [17].
@@
We have applied QMCL to a closure problem for the shallow water equations on a periodic, one dimensional domain. The numerical results demonstrate that QMCL can correctly predict the dynamics of the out of sample initial conditions tested in this work, generating height and momentum fields that agree both qualitatively and quantitatively with the true data. At the same time, it struggles to fully capture the strong spatial variations observed in the subgrid fluxes, leading to more diffusive evolution. Overall, we believe that the results act as a successful proof of concept for the extension of QMCL to spatiotemporal dynamics, especially given the modest amount of data used for its training. Everything else kept the same, it is expected that increasing the amount of training data or the spectral resolution $L$ will lead to improved prediction accuracy.
5.1 Computational aspects
Extending the QMCL framework to spatiotemporal dynamics comes with the challenge of increased computational cost for its implementation, primarily due to the increased size of the training datasets required to capture the main features of dynamics that evolves over both space and time. As such, one key area of focus for future work is reducing the cost of implementation of QMCL and improving its scaling with sampling parameter $N$, spatial grid size $M$ and spectral resolution parameter $L$.
In the offline stage the most expensive operation is computing the leading $L$ eigenfunctions of the $NM\times NM$ symmetric kernel matrix $\bm{K}_{N}$ representing kernel integral operator (8). Even for moderate size parameters $N$ and $M$, solving such an eigenvalue problem can quickly become very expensive. As explained in Section 4.3 to reduce this computational cost when generating our numerical results we performed a partial Cholesky factorization of $\bm{K}_{N}$ to approximate it by a lower rank matrix. The fact that the partial Cholesky factorization preserves the positivity of matrix entries means that we can use the approximate kernel matrix to perform its bistochastic normalization and solve the resulting eigenvalue problem with reduced cost [9, 40].
As an alternative we also tried using the random Fourier features (RFF) method to approximate $\bm{K}_{N}$ by a lower rank matrix [31]. However, unlike the partial Cholesky factorization, the RFF method does not preserve positivity of matrix entries, which means that attempting to normalize the resulting approximate matrix is not a well defined operation. Additionally, we empirically found that a relatively large number of Fourier features was required to obtain a satisfactory approximation of kernel matrix $\bm{K}_{N}$ compared to the rank parameter $r$ required for the partial Cholesky factorization. We intend to continue investigating ways of reducing the cost of computing the kernel eigenfunction basis.
One way of alleviating this issue is to use the RFF method to approximate the out of sample evaluation of kernel (9), thereby eliminating the dependence on the training data [31, 21]. In our case this approach requires keeping in memory one $L\times L$ matrix for each random Fourier feature used. Namely, using $r$ Fourier features requires holding in memory $rL^{2}$ floating point numbers of the employed precision. In our experiments we have found that a large number of Fourier features was required to resolve the fine spatial features of our state vectors, rendering this memory cost requirement unsustainable.
A source of potential numerical instability in the implementation of QMCL is the renormalization of the quantum state after its bayesian conditioning. As briefly discussed in Section 4.4, this can become problematic when no part of the training dataset is identified by the conditioning kernel as being similar to the current resolved state, leading to an approximately zero feature vector and conditioned state. More generally, obtaining an approximately zero conditioned quantum state can happen when the prior quantum state and the observation encoded in the feature vector have disjoint supports. One potential way of addressing this issue is resetting the prior quantum state to an uninformative state and repeating the conditioning operation. This requires erasing the historical information accumulated in the prior state, thereby placing a higher level of trust on the dynamical relevance of the observation. Building a better theoretical understanding of the conditions under which such a scenario might occur and the computational ways in which it can be addressed is a topic of future work.
Finally, the successful application of QMCL relies on the precise calibration of kernel bandwidth $\epsilon$, number of delays $Q$ and conditioning stencil size $J$. Having efficient, reliable algorithms for selecting these parameters based on a training dataset is very important. Tuning bandwidth parameter $\epsilon$ can be achieved reliably using the algorithm developed in [12], although corrections are usually required when working with small training datasets. When applicable, choosing the number of delays $Q$ can be informed by the dimension of the attracting state space region sampled by the dynamical trajectories generating the training data [35]. However, such dimension estimates are rarely available for real world datasets and their underlying dynamics. It would be interesting to consider computational algorithms that can estimate the right range of values for $Q$ based on the given training dataset, potentially also accounting for the presence of noise [45, 26, 6].
5.2 Theoretical aspects
One of the distinguishing features of the QMCL framework is that the classical dynamics is first embedded into the noncommutative setting of operator algebras used in quantum mechanics, and only then discretized to finite dimension. Namely, instead of directly discretizing functions on Hilbert space $H$ to obtain functions on finite dimensional subspace $H^{L}$, we first map functions on $H$ to operators in $B(H)$ and then discretize them to obtain operators in finite dimensional $B(H^{L})$. This process renders the discretization employed by QMCL positivity preserving, with positive classical observables and probability density functions mapped to positive operators [17].
In addition to structure preserving properties, this discretization process embeds the discretized quantities in a space of larger dimension than direct discretization. More specifically, direct discretization maps real functions on $H$ to functions on the $L$-dimensional subspace $H^{L}$. The QMCL discretization maps real functions on $H$ to self adjoint real operators in $B(H^{L})$, which is a subspace of dimension $\frac{1}{2}L(L+1)$. It would be interesting to explore whether this difference in dimension along with the noncommutative nature of $B(H^{L})$ allow QMCL to encode more information about the original dynamics compared to closure schemes performed in $H^{L}$. Using tools from control theory and information theory one should be able to quantify the differences between the two discretization approaches and their consequences for downstream closure schemes.
Figure 1: Space-time heatmaps for trajectory with $\delta=0.25$. Top row is the results for the resolved state variables; bottom row for the subgrid fluxes. For each row, from left to right: the first two plots show the results for the height field $h$, with the true results on the left and the predicted results on the right; the next two plots are the results for the momentum field $q$, with true results on the left and predicted on the right.
Figure 2: Space-time heatmaps for trajectory with $\delta=0.75$. The layout of the plots is as in Figure 1.
Figure 3: Spatial profiles of final snapshot for trajectory with $\delta=0.25$. Top row is the results for resolved state and fluxes for the height variable; bottom row for the momentum variable. In each row, true results are on the left subplot, predicted results on the right subplot. Within each subplot, full lines are used for the resolved state with their values indicated on the left axis; dashed lines are used for the corresponding fluxes with their values indicated on the right axis.
Figure 4: Spatial profiles of final snapshot for trajectory with $\delta=0.75$. The layout of the plots is as in Figure 3.
Appendix A Implementation details
@@
The offline stage consists of the operations needed to build the QMCL scheme using the training data $\{\hat{x}_{n}\}_{n=0}^{N-1}\subset\mathcal{X}$ and $\{z_{n}\}_{n=0}^{N-1}\subset\mathcal{Z}$. This involves the following sequence of steps.
@@
    We use samples $\{\hat{x}_{n}\}_{n=0}^{N-1}$ to compute the leading $L$ eigenfunctions $\{\phi_{\ell}\}_{\ell=0}^{L-1}$ of integral operator (8), defining subspace $H_{\Omega,N}^{L}$ as their linear span. The eigenfunction basis is represented by matrix $\bm{\Phi}\in\mathbb{R}^{NM\times L}$ holding one eigenfunction in each column.
@@
    We use the flux samples $\{z_{n}\}_{n=0}^{N-1}$ to form two quantum observables and project them to operators $A^{(0)}$, $A^{(1)}\in B(H_{\Omega,N}^{L})$ represented by matrices $\bm{A}^{(0)}$, $\bm{A}^{(1)}\in\mathbb{R}^{L\times L}$. The first observable is built using the flux samples for the height $h$ variables; the second using those for the momentum $q$ variables.
@@
    We define the discrete field of quantum states $\rho\in L^{2}(S,\nu_{M};B(H_{\Omega,N}^{L}))$ with one state corresponding to each coarse grid cell. Each quantum state $\rho_{m}=\rho(s_{m})\in Q(H_{\Omega,N}^{L})\subset B(H_{\Omega,N}^{L})$, $s_{m}\in S_{M}$, is represented by a vector $\underline{\rho}_{m}\in\mathbb{R}^{L}$.
@@
In the online stage we use the operators built in the offline stage to timestep the surrogate dynamics. We denote by $\tilde{x}\in\mathcal{X}$ and $\tilde{z}\in\mathcal{Z}$ the resolved state and fluxes predicted by QMCL; this is to distinguish them from the true resolved state $\hat{x}$ and fluxes $z$. The online stage consists of the following steps.
@@
    We initialize the quantum states representing each grid cell using an uninformative initial condition: each quantum state is set to the projection along the unit function $1_{\Omega}\in H_{\Omega,N}^{L}$ that takes the value $1$ everywhere.
@@
    We use the quantum states and quantum observables to compute the discrete field of flux values $\tilde{z}$ corresponding to two flux values for each grid cell.
@@
      We update the resolved state using the computed fluxes.
@@
      We update the quantum states using the transfer operator.
@@
      We compute the updated flux values.
In the timestepping loop as presented above, the quantum states are conditioned at every timestep. In our presented numerical results we have instead elected to condition the states less frequently, by only performing bayesian conditioning every $N_{r}$ timesteps with parameter $N_{r}\in\mathbb{N}$.

### v2

@@ 5.4 Discussion
In the following paragraphs we discuss several aspects of the presented closure scheme and how they inform future research directions.
Computational aspects
Extending the QMCl framework to spatiotemporal dynamics comes with the challenge of increased computational cost for its implementation, primarily due to the increased size of the training datasets required to capture the main features of dynamics that evolves over both space and time. As such, one key area of focus for future work is reducing the cost of the numerical implementation of QMCl and improving its scaling with sampling parameter $N$, spatial grid size $M$ and spectral resolution parameter $L$.
In the offline stage the most expensive operation is computing the leading $L$ eigenfunctions of the $NM\times NM$ symmetric kernel matrix $\bm{K}_{N}$ representing the kernel integral operator (8). Even for moderate size parameters $N$ and $M$, solving such an eigenvalue problem can quickly become very expensive. As explained in Section 5.3, to reduce this computational cost when generating our numerical results we performed a partial Cholesky factorization of $\bm{K}_{N}$ to approximate it by a lower rank matrix using the randomized numerical algorithm developed in [66]. We then used the approximate kernel matrix to perform its bistochastic normalization and solve the resulting eigenvalue problem with reduced cost [67].
As an alternative we also tried using the random Fourier features (RFF) method to approximate $\bm{K}_{N}$ by a lower rank matrix [81]. However, we found that a relatively large number of Fourier features was required to obtain a satisfactory approximation of kernel matrix $\bm{K}_{N}$ compared to the rank parameter $r$ required for the partial Cholesky factorization. The poor approximations obtained using the RFF method led to issues in the downstream bistochastic normalization of the approximate kernel matrix, often encountering negative or zero row sums that prevented its normalization. In addition to our empirical results, the difference in the accuracy of approximations obtained using partial Cholesky and RFF is also supported by the slower convergence rate of the RFF method [66]. We intend to continue investigating effective ways of reducing the cost of computing the kernel eigenfunction basis.
One way of alleviating this issue is to use the RFF method to approximate the out of sample evaluation of kernel (9), thereby eliminating the dependence on the training data [81, 82]. In our case this approach requires keeping in memory one $L\times L$ matrix for each random Fourier feature used. Namely, using $r$ Fourier features requires holding in memory $rL^{2}$ floating point numbers of the employed precision. In our experiments we found that a large number of Fourier features was required to resolve the fine spatial features of the state vectors, rendering this memory cost requirement unsustainable.
@@
In QMCl the surrogate fluxes are computed as the expected value of an observable with respect to a given quantum state. As a result, to compute fluxes that vary strongly over the spatial domain one has to ensure that the used quantum states are very sharp: obtaining a clear maximum in some regions of product space $\Omega$ and being close to zero everywhere else, formally similar to a finite sum of Dirac delta measures. Working with such sharp quantum states poses several numerical challenges.
@@
Producing accurate but very sharp quantum states requires a very precise choice of bandwidth parameter $\epsilon^{\prime}$ used in kernel (9) for the bayesian conditioning. Having such a selective kernel means that the resulting quantum state leads to highly accurate fluxes when the correct part of the training dataset is identified as similar to the current resolved state. However, it can lead to strongly inaccurate fluxes when regions of the training dataset are misidentified as similar, compared to using a less selective kernel that leads to wider averaging of the employed quantum observable.
@@
Moreover, the renormalization of the quantum state after its bayesian conditioning can become a source of numerical instability when working with a very selective kernel. This can occur when no part of the training dataset is identified by the conditioning kernel as being similar to the current resolved state, leading to an approximately zero feature vector and conditioned state. More generally, obtaining an approximately zero conditioned quantum state can happen when the prior quantum state and the observation encoded in the feature vector have disjoint supports. In those cases, attempting to normalize the quantum state to unit norm can yield unpredictable results. Working with a wider, less selective kernel makes this scenario less likely to occur, albeit with the tradeoff of less informative conditioning and thus less accurate predicted fluxes.
Another way of addressing this issue is resetting the prior quantum state to an uninformative state and repeating the conditioning operation. This requires erasing the historical information accumulated in the prior state, thereby placing a higher level of trust on the dynamical relevance of the observation. Building a better theoretical understanding of the conditions under which such a scenario might occur and the computational ways in which it can be addressed is a topic of future work.
@@
The challenges involved in producing and working with a very selective kernel function motivate our use of a kernel bandwidth tuning algorithm [73] and of the introduction of variable bandwidth [76, 64]. These measures help us define a kernel function that is tailored to the characteristics of our training dataset, which in turn allows us to strike a reasonable compromise between numerical stability and sharpness of the kernel used for bayesian conditioning. As a result, we are able to construct a closure scheme that can predict the main features of the considered dynamics with reasonable computational cost and without numerical instabilities.
Finally, the successful application of QMCl relies on the precise calibration of kernel bandwidth $\epsilon$, number of delays $Q$ and conditioning stencil size $J$. Having efficient, reliable algorithms for selecting these parameters based on a training dataset is very important. The tuning of bandwidth parameter $\epsilon$ can be achieved using the algorithm developed in [73] or the median rule studied in [83]. When applicable, choosing the number of delays $Q$ can be informed by the dimension of the attracting state space region sampled by the dynamical trajectories generating the training data [84]. However, such dimension estimates are rarely available for real world datasets and their underlying dynamics. It would be interesting to consider computational algorithms that can estimate a suitable range of values for $Q$ based on the given training dataset, potentially also accounting for the presence of noise [85, 86, 87].
Theoretical aspects
One of the distinguishing features of the QMCl framework is that the classical dynamics is first embedded into the noncommutative setting of operator algebras used in quantum mechanics and only then discretized to finite dimension. Namely, instead of directly discretizing functions on Hilbert space $H$ to obtain functions on finite dimensional subspace $H^{L}$ we first map functions on $H$ to operators in $B(H)$ and then discretize them to obtain operators in finite dimensional $B(H^{L})$. This process renders the discretization employed by QMCl positivity preserving, with positive classical observables and probability density functions mapped to positive operators [55].
In addition to structure preserving properties, this discretization process embeds the discretized quantities in a space of larger dimension than direct discretization. More specifically, direct discretization maps real functions on $H$ to functions on the $L$-dimensional subspace $H^{L}$. The QMCl discretization maps real functions on $H$ to selfadjoint real operators in $B(H^{L})$ which is a subspace of dimension $\frac{1}{2}L(L+1)$. It is interesting to explore whether this difference in dimension along with the noncommutative nature of $B(H^{L})$ allow QMCl to encode more information about the original dynamics compared to closure schemes performed in $H^{L}$. Using tools from control theory [88] and quantum information theory [89] one should be able to quantify the differences between the two discretization approaches and their consequences for downstream closure schemes.
Another question concerns the out of sample performance of QMCl. In Sections 4.5 and 5.3 we considered training data generated by a finite collection of initial conditions without reference to an explicit process of generating these initial conditions. A natural extension of this scenario is to consider an invariant measure $\mu$ admitting a disintegration into conditional measures $\mu_{i}$ sampled from a probability distribution of initial conditions. Interesting questions in this direction are to characterize the behavior of our closure scheme as the number of sampled trajectories increases and derive associated out of sample error bounds.
@@ 6 Conclusion
We have developed a framework for the dynamical closure of spatiotemporal dynamics which combines elements of the quantum mechanical closure (QMCl) framework developed in [56, 57, 55] and the vector valued spectral analysis (VSA) technique of [65]. A key element of the present extended framework is that it encodes statistical information about unresolved degrees of freedom in the form of a field of quantum density operators on the spatial domain. This results in a flexible architecture where density operators are collocated with the discretization cells of the coarse model for the resolved dynamics. The density operators communicate with the resolved dynamics by providing the required flux terms through expectations of quantum observables and by receiving information through operator valued feature maps. Moreover, our approach inherits the symmetry factorization properties of VSA during the basis construction process and can also be used with training data consisting of multiple dynamical trajectories. As with the original QMCl work, the process of embedding classical dynamics into the noncommutative setting of quantum mechanics leads to positivity preserving discretizations of classical observables and probability density functions.
@@
We have applied QMCl to a closure problem for the shallow water equations on a periodic one dimensional domain. The numerical results demonstrate that QMCl can effectively predict the dynamics of the out of sample initial conditions tested in this work, generating height and momentum fields that agree both qualitatively and quantitatively with the true data. As expected from a dynamical system that demonstrates strong spatial variations in the subgrid fluxes, the surrogate dynamics is more diffusive than the true dynamics generated by a high resolution reference model, and eventually struggles to fully capture the sharp spatial features of the true solution.
With that said, we believe that the results act as a successful proof of concept for the extension of QMCl to spatiotemporal dynamics, especially given the modest amount of data used for its training. Everything else kept the same, we expect that increasing the size of the training dataset $N$ or the spectral resolution $L$ will lead to improved prediction accuracy.
Future research directions stemming from this work include improving the computational cost of the data driven implementation of QMCl (particularly the cost of quantum state conditioning in the online stage), as well as the development of strategies for sampling initial conditions in the training stage and kernel tuning/learning. Additionally, we plan to explore applications of QMCl to systems exhibiting spatiotemporal chaos.
Acknowledgements
DG and JS acknowledge support from the U.S. Department of Energy under grant DE-SC0025101. CV was supported as a postdoctoral researcher from this grant. DG acknowledges support from the U.S. Department of Defense, Basic Research Office under Vannevar Bush Faculty Fellowship grant N00014-21-1-2946. DCF was supported as a PhD student from this grant. We thank Ilon Joseph, Michael Montgomery and Ilya Timofeyev for helpful conversations on the topics of this work.
Appendix A Data driven implementation
@@
The offline stage consists of the operations needed to build the QMCl scheme using the training data $\{\hat{x}_{n}\}_{n=0}^{N-1}$ and $\{z(x_{n},\cdot)\}_{n=0}^{N-1}$. This involves the following sequence of steps.
@@
    We use the samples $\{\hat{x}_{n}\}_{n=0}^{N-1}$ to compute the leading $L$ eigenfunctions $\{\phi_{\ell}\}_{\ell=0}^{L-1}$ of integral operator (8), defining subspace $H_{\Omega,N}^{L}$ as their linear span. The eigenfunction basis is represented by matrix $\bm{\Phi}\in\mathbb{R}^{NM\times L}$ holding one eigenfunction in each column.
@@
    We use the flux samples $\{z(x_{n},\cdot)\}_{n=0}^{N-1}$ to form two quantum observables and project them to operators $A_{0}$, $A_{1}\in B(H_{\Omega,N}^{L})$ represented by the matrices $\bm{A}_{0}$, $\bm{A}_{1}\in\mathbb{R}^{L\times L}$. The first observable is built using the flux samples $\{z_{0}(x_{n},\cdot)\}_{n=0}^{N-1}$ corresponding to the height $h$ variables; the second using $\{z_{1}(x_{n},\cdot)\}_{n=0}^{N-1}$ corresponding to the $q$ variables.
@@
    We define the discrete field of quantum states $\rho\in L^{2}(S,\nu_{M};B(H_{\Omega,N}^{L}))$ with one state corresponding to each coarse grid cell. Each quantum state $\rho(s_{m})\in Q(H_{\Omega,N}^{L})\subset B(H_{\Omega,N}^{L})$, $s_{m}\in S_{M}$ is represented by a vector $\underline{\rho}_{m}\in\mathbb{R}^{L}$.
@@
In the online stage we use the operators built in the offline stage to timestep the surrogate dynamics. We denote by $\tilde{x}\in\mathcal{X}$ the resolved state predicted by QMCl to distinguish it from the true resolved state $\hat{x}$. The online stage consists of the following steps.
@@
    We initialize the quantum states $\rho$ representing each grid cell using an uninformative initial condition: each quantum state $\rho(s_{m})$ is set to the projection along the unit function $1_{\Omega}\in H_{\Omega,N}^{L}$ that takes the value 111 everywhere.
@@
    We use the quantum states $\rho$ and quantum observables to compute the surrogate fluxes $\tilde{z}(\rho)$ corresponding to two flux values for each grid cell.
@@
      We update the resolved state $\tilde{x}$ using the surrogate fluxes.
@@
      We update the quantum states $\rho$ using the transfer operator.
@@
      We compute the updated surrogate fluxes $\tilde{z}(\rho)$.
In the timestepping loop as presented above, the quantum states are conditioned at every timestep. In our presented numerical results we have elected to condition the states less frequently, by only performing bayesian conditioning every $N_{r}$ timesteps with parameter $N_{r}\in\mathbb{N}$.